### v1

# HoloDx: Knowledge- and Data-Driven Multimodal Diagnosis of Alzheimer’s Disease

Abstract
Accurate diagnosis of Alzheimer’s disease (AD) requires effectively integrating multimodal data and clinical expertise. However, existing methods often struggle to fully utilize multimodal information and lack structured mechanisms to incorporate dynamic domain knowledge. To address these limitations, we propose HoloDx, a knowledge- and data-driven framework that enhances AD diagnosis by aligning domain knowledge with multimodal clinical data. HoloDx incorporates a knowledge injection module with a knowledge-aware gated cross-attention, allowing the model to dynamically integrate domain-specific insights from both large language models (LLMs) and clinical expertise. Also, a memory injection module with a designed prototypical memory attention enables the model to retain and retrieve subject-specific information, ensuring consistency in decision-making. By jointly leveraging these mechanisms, HoloDx enhances interpretability, improves robustness, and effectively aligns prior knowledge with current subject data. Evaluations on five AD datasets demonstrate that HoloDx outperforms state-of-the-art methods, achieving superior diagnostic accuracy and strong generalization across diverse cohorts. The source code will be released upon publication acceptance.
Multimodal Alzheimer’s Disease Diagnosis, Knowledge- and Data-Driven Framework, Large Language Models, Computer-aided Diagnosis
1 Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that leads to memory loss, cognitive decline, and ultimately, loss of independence. Accurate diagnosis of AD is crucial for implementing appropriate treatment strategies, planning for future care, and providing support to patients and their families. Recent advancements [1, 2] in passive immunotherapies, such as Lecanemab and Donanemab, have shown promise in targeting amyloid-beta plaques, a hallmark of AD pathology. However, diagnosing AD, especially in its early stages, remains challenging due to subtle clinical symptoms and overlap with normal aging processes. Traditional diagnostic approaches, often relying on unimodal data such as neuroimaging or clinical assessments, struggle to capture these nuanced changes [3, 4]. To address these limitations, multimodal learning has gained attention for its ability to integrate diverse data sources, such as neuroimaging, clinical records, and genetic information, offering a more holistic perspective on AD’s early pathological changes [5, 6]. Despite the promise of data-driven approaches, significant challenges remain: 1) Data bias, which focuses on prevalent patterns while often neglecting rare and atypical cases, and 2) A lack of interpretability, as many models function as opaque black boxes, making it difficult to trace causal relationships and ensure reliability. These limitations hinder the consistency, robustness, and trustworthiness of existing computer-aided diagnosis (CAD) systems. Addressing these shortcomings requires a synergistic paradigm, as shown in Fig. 1, which integrates the strengths of multimodal data-driven models with domain knowledge, enhancing interpretability, reliability, and generalizability in the context of AD and its early diagnosis.
To effectively integrate data and domain knowledge, current methods have explored incorporating domain expertise into multimodal diagnostic systems. Many of these methods rely on predefined medical ontologies, such as knowledge graphs, to embed expert insights into the diagnostic process [7]. For example, the Alzheimer’s Disease Knowledge Graph (ADKG) [8] encodes relationships between key concepts using curated medical abstracts. While these methods provide a structured representation of domain knowledge, they face significant limitations. Their static knowledge bases, like ADKG, make it difficult to adapt to individual patient data or changing clinical conditions, restricting their capacity to offer personalized insights. Moreover, despite attempts to compile extensive resources, these graphs typically represent only a portion of the ever-expanding and evolving body of medical knowledge, limiting their comprehensiveness. Also, integrating large-scale but mismatched knowledge could lead to redundancy and inconsistencies, making cross-modality alignment more difficult and compromising diagnostic reliability. These challenges highlight the need for innovative frameworks capable of adaptively incorporating coherent and personalized domain knowledge while maintaining synergy across modalities.
Large language models (LLMs) have emerged as a promising tool for knowledge-driven medical diagnosis, offering expert-like reasoning and vast general knowledge. For example, HuatuoGPT-V [9], a multimodal model designed for medical consultation, demonstrates this potential. By leveraging extensive domain knowledge from sources like PubMed, LLMs can convey general medical facts through natural language interactions across multimodal data types. In addition to broad knowledge, LLMs can provide personalized insights by tailoring general information to individual patient contexts, improving clinical applicability and prediction accuracy. Also, expert knowledge is crucial for more context-specific interpretations of complex cases, filling the gaps left by LLMs’ generalized approach. To combine these complementary sources, we propose merging LLM knowledge with domain expertise to create a unified knowledge bank, ensuring consistency, reducing redundancy, and encoding context-relevant information for more accurate and personalized diagnosis.
Building on the knowledge bank, we propose HoloDx, a novel knowledge- and data-driven CAD framework, as shown in Fig. 2. HoloDx integrates multimodal data sources, such as imaging, cognitive tests, biospecimens, and genetic information, to provide accurate and universally applicable clinical assessments. By combining general insights from LLMs with specialized expertise from domain professionals, the system ensures reliable decision-making. To integrate this knowledge effectively, we design a knowledge injection module based on knowledge-aware gated cross-attention, enriching data samples with domain-specific insights and aligning visual and textual features. Additionally, a memory injection module with the proposed prototypical memory attention dynamically collects prior patient experience, mimicking expert knowledge accumulation and improving adaptability and performance.
Overall, our contributions are summarized as follows:
A novel Dual-Driven CAD framework: We propose HoloDx, a knowledge- and data-driven model designed to improve multimodal diagnosis of Alzheimer’s disease and its early stages. By leveraging dynamic and adaptive domain knowledge learning, HoloDx achieves superior diagnostic accuracy, consistently outperforming state-of-the-art baselines across five datasets.
Injection of domain knowledge and dynamic prototypical memory: We design a knowledge injection module to incorporate dynamic domain knowledge from medical LLMs and experts, ensuring informed multimodal representations. A dynamic prototypical memory mechanism further enables the model to adapt and refine its understanding by leveraging knowledge from past cases for improved performance.
2 Related work
2.1 Multimodal CAD
The integration of diverse modalities of medical data has become a prominent approach in advancing medical diagnosis [10, 11], enabling machine learning models to identify disease patterns more effectively than relying on single modalities. For example, Kim et al. [12] develop a heterogeneous graph learning method to fuse multimodal medical data efficiently. Similarly, Zhou et al. [6] introduce the IRENE model, which employs Transformer-based architectures to combine modality-specific low-level embeddings for diagnostic purposes. Multimodal integration has also succeeded in diagnosing conditions such as breast cancer, where combining imaging and genomic information has improved diagnostic outcomes [13]. Chen et al. [5] propose a transformer-based framework for aligning and fusing multimodal medical data to classify Alzheimer’s disease. These efforts highlight the growing potential of multimodal CAD systems to improve diagnostic precision and reliability. However, these approaches often rely solely on data-driven techniques, which limits their ability to handle complex medical scenarios where domain expertise is crucial. Without integrating domain knowledge, such models may fail to resolve conflicts between modalities or interpret rare and ambiguous cases effectively. This gap underscores the need for combining multimodal data integration with domain knowledge to enhance interpretability, consistency, and diagnostic reliability.
2.2 Knowledge-driven Multimodal CAD
The integration of domain knowledge has become a cornerstone strategy in healthcare to enhance patient phenotyping and clinical prediction accuracy. By encoding medical expertise into computational frameworks, models gain improved capacity to prioritize clinically relevant biomarkers and diagnostic patterns. Notable implementations include KAME [7], which employs attention-weighted medical knowledge graphs to model multi-disease correlations across clinical encounters. Subsequent architectures like LDAM [14], CAML [15], and LERP [16] further advance this paradigm through cross-attention mechanisms that align laboratory findings with disease-specific risk profiles. Recent innovations such as EHR-KnowGen [17] demonstrate enhanced diagnostic capability through multi-granular knowledge fusion, combining fine-grained symptom ontologies with coarse-grained treatment pathways. Despite these advancements, current methodologies exhibit three fundamental constraints: (1) Static knowledge representations (e.g., ICD-based hierarchies) fail to adapt to patient-specific clinical trajectories; (2) Fixed graph structures inadequately model multimodal interactions between neuroimaging biomarkers, genetic markers, and longitudinal cognitive assessments; (3) Conventional attention mechanisms impose artificial sparsity constraints, hindering discovery of nonlinear biomarker interactions critical in complex disorders like Alzheimer’s disease.
Our approach bridges this gap by dynamically enriching the knowledge bank with general medical insights derived from large language models (LLMs) and specific knowledge contributed by domain experts, tailored to the factors associated with each subject. This ensures that multimodal data is augmented with both broad and context-specific representations, enabling our method to provide more precise, interpretable, and robust predictions, effectively addressing key limitations in existing knowledge-driven frameworks.
3 Methodology
Dataset | Access | Factors |
ADNI | Public | Image: Hippocampus, Entorhinal Cortex, Amygdala, Thalamus, Cerebral White Matter, Cerebral Cortex, Lateral Ventricle, Cerebellum White Matter, Cerebellum Cortex, Brain Stem, Choroid Plexus, Caudate, Putamen, Pallidum, Ventricle, Vessel |
Demographic: Age, Gender, Education Year, Hand, Racial, Medical History (Neurologic, Cardiovascular, etc., 17 in total) | ||
Cognitive Test: MMSE, MoCA, LDELTOTAL | ||
Biospecimen: Thyroid Stim Hormone, Vitamin B12, Platelets, RBC, WBC, Serum Glucose, Urea Nitrogen, etc. (37 in total) | ||
Genetic: APOE | ||
AIBL | Public | Image: Same as ADNI |
Demographic: Age, Gender, Medical History (Neurologic, Cardiovascular, etc., 9 in total) | ||
Cognitive Test: MMSE, LIMMTOTAL, LDELTOTAL | ||
Biospecimen: Thyroid Stim Hormone, Vitamin B12, Platelets, RBC, WBC, Urea Nitrogen, etc. (12 in total) | ||
Genetic: APOE | ||
RENJI | Private | Image: Same as ADNI |
Demographic: Age, Gender, Education Year, Height, Weight, BMI | ||
Cognitive Test: MMSE, MoCA, AVLT, GDS | ||
Biospecimen: WBC, RBC, Serum Abeta40, Serum Abeta42, Serum PTAU181, Serum PTAU217, etc. (48 in total) | ||
Genetic: APOE | ||
OASIS | Public | Image: Same as ADNI; others: Gender, Age, Hand, Education Year, Socioeconomic Status, MMSE, CDR |
MIRIAD | Public | Image: Same as ADNI; others: Gender, Age, MMSE |
3.1 Knowledge and Data Collection
3.1.1 Factor Extraction
Unlike existing methods [7, 14], which primarily focus on general textual knowledge, our approach introduces fine-grained domain knowledge tailored to the factor level and extends factor extraction to a multimodal setting. We define factors as key elements or features derived from multimodal data that are critical for diagnosis, as listed in Table 4. These factors are categorized into two main types: imaging factors and non-imaging factors:
Imaging factors (): Derived from brain MRI scans using a pre-trained segmentation model [18], imaging factors include critical brain structures, e.g., the hippocampus, ventricles, cortex, etc. Fine-grained domain knowledge is injected at this level, providing insights into how specific anatomical structures relate to AD diagnosis.
Non-imaging factors (): Extracted from clinical data, non-imaging factors include demographics (e.g., age, gender, education), cognitive test scores (e.g., MMSE, MoCA), biospecimen data (e.g., Thyroid Stim Hormone, Vitamin B12), and genetic information (e.g., APOE). Domain knowledge is incorporated into each non-imaging factor, linking specific biomarkers or cognitive metrics to disease diagnosis and clinical interpretations.
The extracted imaging and non-imaging factors are fused to form a set of multimodal matched factors (), aligning diverse modalities into a unified representation enriched with domain knowledge. By integrating fine-grained insights directly at the factor level, our method enables more precise, interpretable, and clinically relevant predictions.
Factor | Knowledge |
Serum ptau217 | LLM: Serum ptau217 is a biomarker that can be used to aid in the diagnosis of Alzheimer’s disease, as higher levels are associated with increased risk and progression of AD. |
Expert: The increase of p-tau217 in the serum of AD patients is usually early and significant, and its diagnostic power is generally superior to that of Abeta42/40 and ptau181. Therefore, p-tau217 holds high clinical value in the early diagnosis of Alzheimer’s disease. | |
MMSE | LLM: The MMSE score serves as a crucial instrument in evaluating cognitive abilities, helping to detect those potentially afflicted with AD. Typically, scores beneath the threshold of 24 on the MMSE suggest potential cognitive deficits, possibly linked to AD. |
Expert: The MMSE score criteria for suspecting cognitive impairment are as follows: 19 for the illiterate, 22 for those with a primary school background, and 26 for individuals with junior high school education or higher. | |
APOE | LLM: APOE is a genetic marker that increases the risk of developing Alzheimer’s disease, with certain alleles being more strongly associated than others. |
Expert: The APOE e4 allele is the strongest genetic risk factor for sporadic Alzheimer’s disease. Compared to the most common APOE e3 allele, individuals with one APOE e4 allele have about 3.7 times the risk of developing Alzheimer’s disease, and those with two APOE e4 alleles face up to a 12-fold increased risk. |
Factors | Serum Abeta40, Serum Abeta42, Serum PTAU181, Serum PTAU217, Serum Abeta42/40, MMSE, MoCA, AVLT, GDS, BMI, Age, Education, Gender, APOE, Hippocampus, Cerebral Cortex, Amygdala |
3.1.2 Factor Knowledge Generation
Based on the extracted factors, we enhance AD diagnosis by generating domain knowledge tailored to them. Our knowledge generator utilizes the extensive capabilities of LLMs to produce a comprehensive set of general insights. Using a text-based prompting strategy, we query the LLM with inputs specifically designed to elicit detailed and relevant information for each factor. For example, to obtain general knowledge about a factor’s relationship with AD diagnosis, we prompt the LLM with:
Q: Please describe the relationship between {factor} and AD diagnosis.
Here, “factor” refers to a specific factor , as listed in Table 1. Examples include the Mini-Mental State Examination (MMSE) scores collected from cognitive tests. The generated responses offer comprehensive descriptions of each factor and its relevance to diagnosing AD, as presented in the LLM knowledge column of Table 2.
While LLMs provide general insights for each factor, clinical precision often requires nuanced interpretations grounded in medical practice or specialized research findings. To ensure context-specific relevance, we integrate the broad knowledge generated by LLMs with domain expertise contributed by medical professionals. Specifically, AD experts enrich the interpretation of critical factors (see Table 3) in our experiments. Their contributions are sampled in the expert knowledge column of Table 2, showcasing the synergy between LLM-generated insights and expert-provided contextual knowledge.
3.1.3 Knowledge Bank
We first construct an offline knowledge bank to effectively manage the integration of general knowledge generated by LLMs and specific knowledge provided by experts. This bank stores knowledge for pre-identified factors, ensuring the model has immediate access to comprehensive domain insights during training and inference. When new factors arise, the LLM is queried to generate relevant general knowledge, which is then combined with expert-provided insights if available and added to the knowledge bank. By dynamically incorporating new information, the knowledge bank not only ensures the system remains current with evolving data but also enhances the model’s adaptability and scalability, enabling it to handle diverse and expanding diagnostic scenarios with precision.
3.1.4 Multimodal Data Collection
The multimodal inputs consist of imaging data, such as structural MRI scans, and non-imaging data, including demographic information, cognitive test results, biospecimen analyses, and genetic profiles. For non-imaging clinical data, such as the MMSE scores, we argue that their numerical values and categorical information are more semantically meaningful when presented in textual descriptions. To this end, we transform raw data entries into descriptive text representations. For example, an MMSE score of 29 is represented as “The MMSE score is 29”. As illustrated in Fig. 2, all non-imaging data are systematically converted into text representations to serve as inputs to our model.
3.2 Multimodal and Knowledge Feature Extraction
With the factor-specific knowledge generated, we proceed to multimodal feature extraction, where image, text, and knowledge representations are systematically processed to capture meaningful and context-aware features for AD diagnosis.
Unimodal Encoder. We employ unimodal encoders to process both visual inputs and textual inputs . For the visual modality, a 3D Vision Transformer [4] is used to encode the input images into vector representations, generating an image embedding that includes a [CLS] token to represent the global image feature. For the textual modality, we use a Longformer Transformer [19], which maps long text representations into feature embeddings. Similarly, a [CLS] token is appended to the beginning of the text input to summarize the whole sentence representation. Each unimodal encoder consists of transformer blocks comprising self-attention (SA) layers and feed-forward networks (FFN).
Unimodal Decoder. To preserve the fine-grained understanding of image and text modalities, we incorporate a unimodal decoder following each unimodal encoder. This lightweight decoder consists of six transformer blocks with an embedding dimension of 768. Each transformer block includes an SA layer and an FFN. For the image modality, a linear projection layer reconstructs the image representation; similarly, another linear projection layer reconstructs the text representation.
Knowledge Encoder. The textual knowledge set, denoted as , comprises factors that encapsulate the aggregated knowledge, which is encoded using a textual knowledge encoder. For consistency and to enable the extraction of robust features within a unified feature space, the knowledge encoder is designed to share weights with the textual encoder. The resulting knowledge embeddings function as the keys and values with the cross-attention layer of the subsequent knowledge injection module, facilitating the seamless integration of domain knowledge into the model.
3.3 Knowledge Injection Module
Given derived visual embeddings , textual embeddings , and knowledge embeddings , we address the critical challenge of effectively integrating domain knowledge with multimodal data. Our proposed knowledge injection module employs a gated cross-attention mechanism to dynamically fuse clinically relevant information while suppressing noise from irrelevant knowledge elements.
First, we concatenate visual and textual embeddings to construct joint queries, keys, and values:
These multimodal embeddings first undergo self-attention (SA) to model interdependencies across modalities:
where the SA layer enables dynamic reweighting of modality-specific features through learned attention patterns.
The contextualized representations then interact with domain knowledge via the Knowledge-Aware Gated (KAG) mechanism. Using as queries, KAG retrieves and injects relevant knowledge elements from (serving as keys/values):
establishing explicit connections between data-driven patterns and clinical knowledge bases.
The full layer transformation integrates these components through residual learning:
As detailed in Fig. 2, the KAG operator implements parametric feature gating:
where and are learnable parameters, denotes the sigmoid activation, and CA represents multi-head cross-attention.
The final patient representation aggregates modality-specific information through concatenation of [CLS] tokens:
where , , and correspond to the [CLS] tokens of visual, textual, and knowledge embeddings respectively.
3.4 Memory Injection Module
In clinical practice, physicians manage diverse patient populations with unique medical histories and clinical profiles. While accumulated experience enhances expertise, conventional methods often struggle to systematically harness insights from growing patient datasets. To address this gap, we propose an innovative memory injection module integrated into our framework, designed as a dynamic memory repository. This module continuously adapts to evolving clinical contexts through collaborative interaction with knowledge injection module, enabling efficient storage and retrieval of insights from historical patient data.
To achieve this, we extend Barraco et al.’s prototypical memory framework for image captioning [20], adapting their memory vector architecture to develop our prototypical memory (PMM) attention mechanism. For a sequential input of mini-batches containing training instance embeddings, we implement dual memory banks and at each network layer. These banks systematically archive keys and values from historical training samples, preserving information within a sliding temporal window spanning iterations. This design captures the evolving geometric structure of key-value relationships across training epochs. The active memory components at the -th iteration are selected as:
where and represent complete key/value sets generated when processing mini-batch through the network layer.
Prototype key vectors are constructed by clustering the key manifold and extracting representative centroids. Formally, for a predefined memory capacity :
where computes cluster centroids from the key memory bank. Value prototypes are subsequently derived through distance-weighted interpolation of values associated with geometrically proximate keys in the manifold. For each key prototype , we compute its corresponding value prototype by aggregating the top- nearest neighbors in the key space:
where retrieves the closest (key, value) pairs to prototype based on distance metric , ensuring contextual alignment between key clusters and their associated value representations.
To maintain memory bank relevance during training while minimizing computational overhead, the memory banks retain a fixed capacity of historical batches, with periodic refresh cycles that selectively incorporate recent key/value streams from network layers. Prototype vectors are regenerated from these updated banks and stored in and . In implementation, this refresh process executes biweekly per epoch, preserving approximately two epochs’ worth of contextual data. The overlapping memory contents between consecutive updates enhance training stability through gradual knowledge transition.
The forward propagation through layers and integrates layer normalization (LN) and residual pathways as:
where PMM attention synergizes memory-derived and input-specific patterns through:
with denoting standard multi-head attention.
3.5 Learning Objectives
3.5.1 Alignment
We employ the image-text contrastive (ITC) loss [23] to align image features and text features generated by the image and text encoders. The ITC loss maximizes similarity for positive image-text pairs while suppressing negative pairs, implemented through normalized cross-entropy over all pairwise similarities. To enhance consistency between input data and domain knowledge in disease diagnosis, we propose the Knowledge-Data Contrastive loss (KDC). This mechanism maintains two dynamically updated queues: a Data queue storing concatenated multimodal features , and a Knowledge queue containing domain embeddings generated by a dedicated knowledge encoder. The KDC loss follows the contrastive paradigm of , pulling matched knowledge-data pairs closer while distancing mismatched pairs. Momentum encoders [24] continuously refresh both queues to maintain feature relevance.
3.5.2 Restoration
Our restorative learning module is designed to enhance the global semantic understanding by incorporating fine-grained visual and textual information. That is, the feature extraction is augmented by a reconstruction learning branch, which includes an image decoder to reconstruct the original image from the representation and minimizes the pixel-level distance between the original image and the reconstructed image : , where presents the distance function that measures similarity between and , e.g., Mean Square Error (MSE), or L1 norm. We use MSE following the common setting [25]. For the textual component, we apply a similar approach. A text decoder is trained to minimize the token-level distance between the original text and the reconstructed text : , where is the distance function measuring text similarity, such as the commonly-used cross-entropy loss.
3.5.3 Classification
The diagnostic classification prediction module leverages each subject’s ground-truth (GT) labels for supervision. Let the disease prediction for the -th subject be denoted as . Given the GT label , the cross-entropy loss for disease prediction is defined as: , where is the total number of training subjects.
The total loss includes the following five terms:
where , , and are constants balancing these terms.
Dataset | Type | #Subjects(F/M) | #MRIs | Age | MMSE | MoCA |
ADNI | NC | 724(407/317) | 3205 | 72.96.3 | 29.01.2 | 26.02.6 |
MCI | 843(344/499) | 4115 | 72.97.4 | 27.52.2 | 23.53.2 | |
AD | 519(222/297) | 2019 | 74.37.3 | 21.64.7 | 16.65.2 | |
AIBL | NC | 361(159/202) | 708 | 73.86.6 | 28.81.7 | - |
MCI | 91(48/43) | 149 | 76.27.0 | 27.12.0 | - | |
AD | 81(31/50) | 140 | 75.28.1 | 19.86.5 | - | |
OASIS | NC | 72(50/22) | 172 | 76.98.1 | 29.20.9 | - |
AD | 64(28/36) | 135 | 76.26.9 | 24.44.5 | - | |
MIRIAD | NC | 30(14/16) | 42 | 68.11.2 | 29.51.0 | - |
AD | 179(104/75) | 244 | 69.97.0 | 17.54.8 | - | |
RENJI | NC | 101(54/47) | 101 | 67.97.9 | 28.31.8 | 22.75.4 |
MCI | 90(57/33) | 90 | 72.29.3 | 23.34.5 | 15.05.5 | |
AD | 180(106/74) | 180 | 71.97.6 | 19.26.0 | 13.85.4 |
Datasets | Methods | NC vs. CI | NC vs. ECI | ||||||
ACC(%) | AUC(%) | SEN(%) | SPE(%) | ACC(%) | AUC(%) | SEN(%) | SPE(%) | ||
ADNI | 3D ResNet[26] | 72.0014.85 | 52.638.86 | 9.7112.20 | 95.568.89 | 67.697.53 | 59.6111.04 | 32.3324.58 | 86.8913.01 |
3D ViT [4] | 76.0016.11 | 58.9115.58 | 25.0429.52 | 92.788.89 | 63.0813.23 | 54.5714.01 | 30.6724.73 | 78.4814.87 | |
BERT [27] | 85.337.77 | 78.258.68 | 64.7618.65 | 91.744.42 | 81.539.23 | 80.2911.06 | 84.3316.38 | 86.2515.33 | |
RoBerta [28] | 86.679.43 | 81.589.71 | 70.1018.67 | 93.069.04 | 87.697.84 | 84.3810.45 | 83.0019.04 | 85.7712.18 | |
Longformer[19] | 88.008.84 | 86.5410.04 | 79.6117.03 | 93.469.70 | 86.155.75 | 85.837.58 | 80.3316.75 | 81.3311.26 | |
GIT [29] | 84.703.43 | 54.624.67 | 98.571.75 | 10.678.79 | 84.114.78 | 71.055.24 | 83.125.33 | 52.0011.27 | |
IRENE [6] | 87.064.78 | 78.447.36 | 90.885.34 | 66.0014.20 | 86.472.99 | 78.106.41 | 90.194.18 | 65.9914.20 | |
Alifuse [5] | 87.175.38 | 82.468.15 | 92.833.78 | 72.0913.22 | 87.054.77 | 78.447.39 | 90.885.34 | 73.4411.20 | |
HoloDx (ours) | 93.335.96 | 91.837.64 | 88.6715.72 | 95.009.99 | 92.822.99 | 90.095.01 | 96.602.11 | 83.599.74 | |
AIBL | 3D ResNet[26] | 65.887.58 | 55.179.12 | 70.687.97 | 39.6714.31 | 60.598.84 | 58.4514.49 | 61.568.09 | 55.3326.72 |
3D ViT [4] | 67.185.23 | 55.757.89 | 82.076.65 | 29.4215.61 | 66.153.77 | 54.467.20 | 81.315.46 | 27.6016.27 | |
BERT [27] | 86.672.51 | 86.622.13 | 86.574.18 | 86.654.77 | 85.294.16 | 71.774.99 | 91.554.90 | 52.0011.27 | |
RoBerta [28] | 87.694.10 | 86.355.48 | 88.484.91 | 84.2310.38 | 88.824.32 | 82.074.97 | 91.485.82 | 72.6712.89 | |
Longformer[19] | 88.213.84 | 86.244.97 | 90.723.05 | 81.769.71 | 88.234.16 | 75.638.57 | 93.604.75 | 57.6719.60 | |
GIT [29] | 84.355.43 | 82.625.13 | 88.575.18 | 84.654.77 | 83.293.16 | 82.774.79 | 89.555.90 | 64.2012.76 | |
IRENE [6] | 84.106.36 | 71.708.52 | 93.202.16 | 44.1916.55 | 84.295.22 | 81.538.07 | 85.556.20 | 67.0013.24 | |
Alifuse [5] | 86.674.41 | 83.886.27 | 90.013.73 | 77.749.73 | 86.155.28 | 77.997.56 | 87.317.08 | 64.3312.45 | |
HoloDx (ours) | 92.303.24 | 89.385.32 | 96.602.11 | 82.1610.83 | 91.283.07 | 87.954.93 | 94.922.99 | 80.9710.80 | |
RENJI | 3D ResNet[26] | 62.676.79 | 53.994.44 | 15.67 8.41 | 92.3210.70 | 61.336.53 | 53.019.33 | 20.3817.03 | 85.649.79 |
3D ViT [4] | 69.3313.06 | 64.2411.74 | 47.1019.03 | 81.3916.00 | 64.006.80 | 54.179.45 | 16.3814.04 | 91.967.52 | |
BERT [27] | 86.6711.16 | 86.0711.49 | 82.6214.43 | 89.5212.20 | 82.6713.06 | 82.0714.85 | 80.4819.54 | 83.6612.19 | |
RoBerta [28] | 89.336.80 | 87.207.21 | 78.6211.55 | 95.896.33 | 88.004.98 | 89.655.68 | 94.2911.43 | 85.025.17 | |
Longformer[19] | 88.019.79 | 86.2112.25 | 78.9612.23 | 93.469.70 | 88.007.78 | 88.127.42 | 86.9511.34 | 89.2912.20 | |
GIT [29] | 81.3314.24 | 76.4814.36 | 59.6217.24 | 93.3313.33 | 80.0011.93 | 83.2510.31 | 90.4813.13 | 76.0216.28 | |
IRENE [6] | 82.676.80 | 85.065.05 | 90.2912.20 | 79.8413.60 | 84.009.98 | 86.0710.08 | 90.4813.13 | 81.6611.80 | |
Alifuse [5] | 89.337.98 | 88.649.48 | 85.8114.66 | 91.469.16 | 86.6712.65 | 85.4413.48 | 80.1018.62 | 90.789.20 | |
HoloDx (ours) | 94.674.99 | 94.805.17 | 93.817.62 | 95.785.18 | 93.335.96 | 93.376.35 | 90.9511.70 | 95.785.18 |
Methods | NC vs. CI (RENJI) | NC vs. CI (OASIS) | NC vs. CI (MIRIAD) | |||||||||
ACC(%) | AUC(%) | SEN(%) | SPE(%) | ACC(%) | AUC(%) | SEN(%) | SPE(%) | ACC(%) | AUC(%) | SEN(%) | SPE(%) | |
3D ResNet[26] | 62.00 | 52.63 | 9.71 | 95.56 | 55.08 | 49.73 | 5.21 | 94.26 | 55.63 | 50.97 | 12.11 | 89.83 |
3D ViT [4] | 65.33 | 54.00 | 10.00 | 98.01 | 64.46 | 50.58 | 8.83 | 89.33 | 65.39 | 53.60 | 13.08 | 94.12 |
BERT [27] | 82.67 | 85.15 | 90.29 | 80.02 | 65.85 | 68.52 | 46.15 | 90.88 | 87.01 | 59.04 | 19.28 | 98.81 |
RoBerta [28] | 84.00 | 80.54 | 67.76 | 93.33 | 48.70 | 53.27 | 14.68 | 91.86 | 75.79 | 69.52 | 60.39 | 78.65 |
Longformer[19] | 88.00 | 86.52 | 82.48 | 90.56 | 47.48 | 52.89 | 7.27 | 98.51 | 88.07 | 59.64 | 19.28 | 99.97 |
GIT [29] | 77.33 | 69.97 | 59.14 | 80.79 | 68.21 | 66.62 | 53.17 | 80.07 | 75.04 | 71.41 | 58.61 | 84.22 |
IRENE [6] | 69.33 | 69.10 | 64.90 | 73.30 | 73.81 | 71.10 | 61.56 | 80.64 | 74.82 | 72.12 | 62.62 | 81.61 |
Alifuse [5] | 80.00 | 83.97 | 97.14 | 70.80 | 75.18 | 71.92 | 60.53 | 83.32 | 76.19 | 74.91 | 70.43 | 79.39 |
HoloDx (ours) | 90.67 | 88.68 | 81.80 | 95.56 | 78.67 | 74.82 | 60.43 | 89.24 | 89.33 | 88.01 | 81.48 | 78.03 |
4 Experiments and Discussions
4.1 Datasets and Settings
We evaluate our model on four public datasets (ADNI [30], AIBL [31], OASIS [32], MIRIAD [33]) and one private dataset RENJI. The RENJI is derived from the Multimodal Imaging Biomarkers in Degenerative Dementia Cohort (MIMI, ClinicalTrials number: NCT06534658[34]) at Renji Hospital, which is affiliated with the School of Medicine at the Shanghai Jiao Tong University. Table 4 presents the number of subjects and MRIs, as well as the distribution of demographic information, i.e., gender and age, and two key cognitive test values, i.e., MMSE and MoCA, across NC, MCI, and AD categories. Other non-imaging information is included in Table 1.
In accordance with established clinical standards [35, 36, 37], we validate our model on two classification tasks. The first task distinguishes between NC and Cognitive Impairment (CI), where CI includes both MCI and AD. The second task differentiates between NC and Early Cognitive Impairment (ECI), with ECI being used synonymously with MCI, a crucial stage for the preclinical identification of AD. We split the ADNI, AIBL, and RENJI datasets subject-wise into training, validation, and test sets with proportions of 70%, 10%, and 20%, respectively. Also, we perform zero-shot tests on OASIS and MIRIAD datasets to assess inter-center generalization.
Baselines. To evaluate the performance of our model, we compare it with three groups of baselines: (1) Image-only group. This group includes two recent methods commonly used in image classification: 3D ResNet [26], and 3D ViT [4]. (2) Text-only group. We select three prominent text-only baselines: BERT [27], Roberta [28], and Longformer [19]. (3) Multimodal group. This group includes three recent transformer-based models that fuse multimodal information for classification: GIT [29], IRENE [6], and Alifuse [5].
Implementation details For pre-processing, each image is resized and cropped to , with a spacing of [1.0,1.0,1.0] and an image patch size of . The maximum input text length of textual input is set to 512. Following common settings in [25], we set all hidden feature sizes to 768 and the number of heads in multi-head attention (MHA) to 12. The number of layers in the unimodal encoder is set to 12, while the number of layers in the decoder is set to 6. The number of layers in the knowledge and the memory modules is set to 6. The iteration T is 100. The values for are empirically set to 1, 1, and 1, respectively. Our model is trained on four NVIDIA RTX 3090 GPUs with a batch size of , using the AdamW optimizer with a learning rate of 2e-5. Evaluation metrics include diagnosis accuracy (ACC), Area Under Curve (AUC), sensitivity (SEN), and specificity (SPE).
4.2 Experimental Results
4.2.1 Intra-Diagnostic Classification
Table 5 presents the results of our method on three datasets, i.e., ADNI, AIBL, and RENJI, comparing it with eight baselines. Our model outperforms image-only, text-only, and other multimodal models, underscoring the importance of integrating knowledge with imaging and textual data for medical diagnostics. Specifically, our results achieve accuracy and AUC values greater than 90%, except for AIBL’s AUCs, which remain above 88%. Notably, the text-only baselines outperform image-only models on the three datasets, as the text contains more comprehensive and directly relevant diagnostic information. Also, in multimodal approaches, the performance of models like IRENE and Alifuse is limited by challenges posed by long text inputs, resulting in performance not exceeding that of text-only unimodal models with a token length receptive field of 512.
4.2.2 Evaluation on Inter-Center Generalization
To assess the generalization of our method, we conduct inter-center testing on the RENJI, OASIS, and MIRIAD datasets, with ADNI for training. Table 6 presents the performance of our HoloDx compared to baselines. Our method achieves the best performance among all models. However, the testing accuracy on OASIS is relatively lower, at 78.67%, with an AUC of 74.82%, partially due to the limited availability of non-imaging data and the model’s strong reliance on MMSE scores (see the follow-up factor analysis). As shown in Table 4, the average MMSE score for AD in OASIS is relatively high compared to other datasets, making it more challenging to distinguish NC and AD. In contrast, HoloDx performs better on MIRIAD, as its MMSE distribution separates NC and AD more clearly.
4.2.3 Ablation Studies
We conduct three experiments to evaluate the contributions of each component in our framework.
(1) Impact of loss terms. We evaluate the contribution of each loss, i.e., the classification loss , two contrastive losses and , and the combined restorative loss for images and texts. Incorporating all loss components is optimal for maximizing performance across all metrics, as reported in Table 7(a). Using only , the model performance experiences a reduction of 2.04% in ACC and 2.01% in AUC compared to the best-performing configuration. Adding and alongside improves performance, increasing ACC by 1.53% and AUC by 1.43%. This improvement highlights the benefits of incorporating contrastive learning between images and texts. The best performance is achieved by adding , reflecting the importance of fine-grained reconstruction in improving unimodal features extraction.
(2) Impact of model components. To assess the effectiveness of each proposed module, we conduct ablation experiments and report the results in Table 7(b). The configuration labeled “w/o KL, MEM” excludes both the knowledge (KL) and memory (MEM) injection modules, relying solely on visual and textual representations for predictions. The “w/o MEM” configuration removes only the memory module with PMM attention. The reported results confirm the critical role of both modules in AD diagnosis. Also, the use of our proposed KAG cross-attention further improves our performance.
(3) Impact of expert knowledge. The final ablation study examines the impact of expert knowledge, with the results presented in Table 8. This study evaluates the model’s performance with and without the integration of expert-provided knowledge. The findings demonstrate that incorporating expert knowledge significantly enhances diagnostic accuracy and AUC, underscoring its value in complementing the knowledge derived from LLM-generated general information.
Task | ACC(%) | AUC(%) | SEN(%) | SPE(%) |
(a) Loss Terms | ||||
90.264.41 | 87.376.05 | 94.422.61 | 80.3311.49 | |
90.753.02 | 87.422.33 | 95.482.21 | 79.388.78 | |
91.793.77 | 88.804.91 | 95.834.11 | 81.769.71 | |
HoloDx (ours) | 92.303.24 | 89.385.32 | 96.602.11 | 82.1610.83 |
(b) Module Terms | ||||
w/o KL, MEM | 89.233.40 | 86.235.79 | 92.930.49 | 79.549.48 |
w/o MEM | 91.283.07 | 87.944.94 | 94.923.00 | 80.9710.80 |
w/o KAG | 91.792.99 | 88.542.82 | 97.832.87 | 77.249.44 |
HoloDx (ours) | 92.303.24 | 89.385.32 | 96.602.11 | 82.1610.83 |
Type | ACC(%) | AUC(%) | SEN(%) | SPE(%) |
LLM | 91.996.53 | 90.817.54 | 83.6114.04 | 98.013.99 |
Expert | 92.024.98 | 92.644.41 | 96.007.99 | 89.299.97 |
LLM+Expert | 94.674.99 | 94.805.17 | 93.817.62 | 95.785.18 |
4.2.4 Factor Analysis
Shapley analysis [38] is performed on the ADNI, AIBL, and RENJI test sets to identify the numerical features that most significantly influenced the model’s diagnostic decisions (Fig. 3). Across all three datasets, the MMSE score consistently ranks among the top two most influential features. In the AIBL dataset, the Immediate Recall score (LIMMTOTAL), a subcomponent of MMSE, is ranked as the most important feature, and it uniquely appears in this dataset. For both ADNI and AIBL, MMSE, Thyroid Stimulating Hormone, Vitamin B12 levels, and the presence of APOE4 alleles are selected consistently among the top ten factors. These findings align with clinical studies that emphasize the strong association of MMSE scores and other key biomarkers with cognitive impairment and AD diagnosis.
In the RENJI dataset, Auditory Verbal Learning Test (AVLT) scores and its subscore emerge in the top ten factors, closely aligned with the knowledge provided by physicians and clinical studies [39]. AVLT is a test designed to assess auditory memory, which is valuable for detecting early-stage cognitive decline. Another interesting finding is the appearance of BMI and weight in the top ten, consistent with other clinical studies linking these factors to cognitive impairment [40, 41].
4.2.5 Biomarker-Level Validation
To improve clinical accessibility, we use easily available factors that do not involve pathological AD biomarkers, and conduct biomarker-level verification to validate model effectiveness. As shown in the raincloud plots in Fig. 4, our predictions correlate significantly with PET biomarkers (AV45, FDG, PIB) and CSF biomarkers (PTAU, TAU, Abeta) in the ADNI cohort, as well as with PET AV45 and Serum PTAU217 biomarkers in the RENJI cohort, demonstrating statistically significant differences between NC and CI groups (p-values 0.001). Notably, PTAU217 exhibits greater diagnostic efficacy than Serum PTAU181 and Abeta42, further supporting its role in detecting cerebral amyloidosis and predicting cognitive decline [42].
5 Conclusion and Discussion
In this paper, we propose a knowledge- and data-driven method for AD diagnosis, integrating expert-provided and LLM-generated knowledge with multimodal imaging and textual data. At the core of our framework are two key modules: (1) the knowledge injection module, which incorporates domain-specific knowledge from experts and LLM-generated insights, enriching the model’s understanding of AD; and (2) the memory injection module, which stores and retrieves relevant past patient information, enabling the model to leverage historical knowledge for more informed decision-making. To further enhance model performance, we employ contrastive and restorative learning strategies to effectively align and fuse multimodal data with knowledge.
Our experiments on multiple AD datasets show that the proposed method consistently outperforms recent baselines, achieving superior classification performance. Also, we visualize the final diagnosis through Shapley analysis and raincloud plots. Shapley values allow us to identify the most important non-imaging contributions, shedding light on how cognitive and genetic factors influence the model’s predictions. Raincloud plots further demonstrate the alignment between model-predicted diagnoses and clinical biomarkers, ensuring that the model’s decisions are consistent with established diagnostic criteria. In future work, we plan to improve the generalization and cross-domain applicability of our model, ensuring it performs effectively across different datasets. Additionally, we will focus on expanding and updating the knowledge base, integrating more expert-provided and LLM-generated information to enhance the model’s diagnostic capabilities.